OpenStudy (anonymous):

The perimeter of a retangler is 194 inches. The length exceeds the width by 29 inches. Find the width and the lenght please.

OpenStudy (anonymous):

Perimeter of sum of all lengths and widths

OpenStudy (anonymous):

Denote lenght L ,width w

OpenStudy (anonymous):

2L+2W=194

OpenStudy (anonymous):

L-29=W

OpenStudy (anonymous):

Solve these two equation

OpenStudy (anonymous):

I will give medal for just the answer so I can get on with other few problem that I am needed help with and I will give you medal for those answers to want help???

OpenStudy (anonymous):

let width=x let length =x+29 so zer are 2 length n 2 width 2(x)+2(x+29)=194 2x+2x+58=194 x=34 that is width=34 now length=29+34=63

OpenStudy (anonymous):

did zat help??